Find the ordered quintuplet (a,b,c,d,e) that satisfies the system of equations

a+2b+3c+4d+5e=41
2a+3b+4c+5d = 34
3a+4b+5c =15
4a+5b = 16
5a = 20

Answer by math_tutor2020(3824)


Start at the bottom and work your way up.

5a = 20 means a = 4
4a+5b = 16 leads to 4*4+5b = 16 which yields b = 0
3a+4b+5c = 15 leads to 3*4+4*0+5c = 15 which yields c = 3/5
And so on.

The solution you should end up at is (a, b, c, d, e) = (4, 0, 3/5, 118/25, 408/125)

3/5 = 0.6
118/25 = 4.72
408/125 = 3.264
Each decimal value is exact and hasn't been rounded.